OpenStudy (anonymous):

i need a math problem using addition subtraction multiplication or division of radicals that simplifies to 5x...

OpenStudy (swissgirl):

\( \large\frac {(3\sqrt{5x}-\sqrt{5x})}{2}*\sqrt{5x}\)

OpenStudy (swissgirl):

Hey so I want you to check my work and simplify this expression to make sure that it simplifies to 5x CUZ IT MAY NOT :P

OpenStudy (anonymous):

okay no problem